OpenStudy (mony01):

can someone tell me the answer of this or put it into wolfram alpha please!!!!!!

OpenStudy (mony01):

\[\int\limits_{0}^{4}\pi(1-(\frac{ y }{ 4})^{4})^{2}-(1-\frac{ y }{ 4})^{2}dy\]

OpenStudy (whpalmer4):

http://www.wolframalpha.com/input/?i=%E2%88%AB%CF%80(1-(y%2f4)%5E4)%5E2%20-%20(1-(y%2f4))%5E2%20dy Verify that I got the integral entered correctly!

OpenStudy (mony01):

yes it is correctly but my integral is definite going from 0 to 4

zepdrix (zepdrix):

If you type stuff in words, wolfram can usually interpret it pretty well: http://www.wolframalpha.com/input/?i=integral+of+pi*%28%281-%28y%2F4%29%5E4%29%5E2-%281-y%2F4%29%5E2%29+dy+from+y%3D0+to+y%3D4

zepdrix (zepdrix):

Oh the pi isn't only times the first term? crap lemme fix that :\

zepdrix (zepdrix):

`is` only times the first term* blah im just all messed up today -_-